OpenStudy (anonymous):

values of x that are roots of the polynomial. x^2-4x+6.

OpenStudy (cruffo):

Does the quadratic factor?

OpenStudy (anonymous):

yes I think so

OpenStudy (anonymous):

it'll have complex roots

OpenStudy (anonymous):

like square roots? I figured. Which is why I need help

OpenStudy (cruffo):

So, you'll want to use the quadratic formula.

OpenStudy (anonymous):

I know but I still cant figure it out

OpenStudy (cruffo):

sorry, browser crashed...

OpenStudy (anonymous):

its ok

OpenStudy (cruffo):

so , a = 1, b = -4, c = 6, \[x = \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\] \[x = \frac{-(-4) \pm \sqrt{(-4)^2 -4(1)(6)}}{2(1)}\]

OpenStudy (cruffo):

a little simplifying, \[x = \frac{4 \pm \sqrt{16 -24}}{2}\] \[x = \frac{4 \pm \sqrt{-8}}{2}\]

OpenStudy (cruffo):

then \[\sqrt{-8} = i\sqrt{8} = i\sqrt{4 \cdot 2} = 2i\sqrt{2}\]

OpenStudy (cruffo):

So simplifying some more, \[x = \frac{4 \pm 2i\sqrt{2}}{2}\] Now we can factor 2 out of the numerator, and cancel it with the 2 in the denominator, \[x = \frac{2(2 \pm i\sqrt{2})}{2}\] \[x = \frac{2 \pm i\sqrt{2}}{1}\] \[x = 2 \pm i\sqrt{2}\]